16. Suppose U=min(X,Y) and the price of X is 1, the price of Y is 1 and income is $12. If the price of X increases to 2, the income effect (in terms of units of X bought) is

a.

2

b.

-1

c.

0

d.

-2

ANSWER:  

d

POINTS:  

1

17. Suppose U=Min(X,Y) and the price of X is 1, the price of Y is 1 and income is $12. If the price of X increases to 2, the substitution effect is

a.

2

b.

-1

c.

0

d.

-2

ANSWER:  

c

POINTS:  

1
Answer #1

UlX, Y) min(x, y) Budget constraint PxX+ PyY I U(X,Y)= min (X,Y) Y = x put y=x in Budget constraint PxX+ PYY PxX + Pyx X ( Pxx (P², Py, I) ۶۱۱ = 12 2+1 ฯ 12 3 By+Py 4 x (Px, Py, I) IE x (Px, Py, I) 4 - 6 = -2 17) X (Px, Py, I) SE = x (Px, py,